OpenStudy (anonymous):

trig identity: cos^2x-sin^2x/1-tan^2x=cosx

OpenStudy (anonymous):



OpenStudy (anonymous):

Whats your question

OpenStudy (anonymous):

For a proof?

OpenStudy (anonymous):

just to prove it

OpenStudy (anonymous):

I literally ddrw the whole thing out

OpenStudy (anonymous):

and this dang site wont let me show it

OpenStudy (anonymous):

Put sin 2(x + cos 2(x in place of 1

OpenStudy (anonymous):

and cancel out stuff

OpenStudy (anonymous):

so when i cancel everything ill be left with 1/ -tan^2 ?

OpenStudy (anonymous):

I had -1 / -tan^2 and that came out to 1/tan^2

OpenStudy (anonymous):

Give me a second to check if im right

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Nope im wrong! sorry

OpenStudy (noelgreco):

There's a problem with the numerator. You'll never get sin^2 and cos^2 to have different signs in a Pythagorean identity.

OpenStudy (anonymous):

its okay at least u tried

OpenStudy (noelgreco):

The numerator is = cos(2x), but that's not going to help.

OpenStudy (anonymous):

?? it says in the paper cos^2x-sin^2x

OpenStudy (xapproachesinfinity):

is the denominator and \(\huge \rm 1-tan^2x\)

OpenStudy (anonymous):

yeah that is the denominator

OpenStudy (xapproachesinfinity):

we have \(\huge 1-tan^2x=\frac{cos^2x-sin^2x}{cos^2x}\) you can see that \(\huge \rm cos^2x-sin^2x \) will cancel and you are left with cos^2x

OpenStudy (anonymous):

so u switched the denominator?

OpenStudy (noelgreco):

Nice! I was trying difference of squares and Pythagorean identities.

OpenStudy (xapproachesinfinity):

\(\huge \frac{\cancel{cos^2x-sin^2x}}{\frac{\cancel {cos^2x-sin^2x}}{cos^2x}}=\frac{1}{\frac{1}{cos^2x}}\)

OpenStudy (xapproachesinfinity):

flip it over and you get cos^2x

OpenStudy (anonymous):

THANKS got it

OpenStudy (xapproachesinfinity):

oh Paythagoras will work but will take to a longer road

OpenStudy (xapproachesinfinity):

np!